The following information is taken from the records of Erie Corp. (in thousands) for the year ended on December 31:

                                                             2019                              2018


Sales                                                    $1,397                        $1,122

Cost of goods sold                                 935                              814

selling expenses                                    154                               121

general expenses                                    88                                 77

other revenue                                             4                                  7

interest expense                                        2                                   9

income taxes                                           134                                 66
Answer #1

In the vertical analysis each line item will be shown as a percentage of sales in the income statement. So sales will be taken as 100% and other items will be expressed as a percentage of sales.

In 2019,

Sales = $1397 = 100%

Cost of the goods sold = Cost of the goods sold / sales = $935 / $1397 *100 = 66.92%

Selling expenses = Selling expenses / Sales = $154 / $1397 *100 = 11.025%

General expenses = General expenses / sales = $88 / $1397 *100 = 6.299%

Other revenue = Other revenue / sales = $4 / $1397 * 100 = 0.286%

Interest expense = Interest expense / Sales = $2 / $1397 *100 = 0.1431%

Income tax = Income tax / sales = $134 / $1397 *100 = 9.59%

In 2018,

Sales = $1122 = 100%

Cost of the goods sold = Cost of the goods sold / sales = $814 / $1122 *100 = 72.54%

Selling expenses = Selling expenses / Sales = $121 / $1122 *100 = 10.78%

General expenses = General expenses / sales = $77 / $1122 *100 = 6.86%

Other revenue = Other revenue / sales = $7 / $1122 * 100 = 0.6238%

Interest expense = Interest expense / Sales = $9 / $1122 *100 = 0.8021%

Income tax = Income tax / sales = $66 / $1122 *100 = 5.88%

Changes from 2018 to 2019

Sales : Sales increased in 2019 as compared to 2018, so it is a favourable change.

Cost of the goods sold: In 2018, it was 72.54% and in 2019 it is 66.92%. It means cost of the good sold percentage has decreased, which is a favorable change, because cost is decreased.

Selling expenses: In 2018, it was 10.78% and in 2019 it is 11.25%. It means that expenses has increased so it is a unfavorable change.

General expenses : , it was 6.86% and in 2019 it is 6.299%. It means that the expenses have decreased so it is a favorable change.

Other revenue: In 2018, it was 0.6238% and in 2019 it is 0.286%. It means that revenue as a percentage of sales is declined which is a unfavorable change.

Interest expenses : , it was 0.8021% and in 2019 it is 0.1421%. It means that the expenses have decreased so it is a favorable change.

Income tax: it was 5.88% and in 2019 it is 9.59%. It means that income tax has increased which is a unfavorable change.